Gucchi:

An inequality is shown: −np − 4 ≤ 2(c − 3) Which of the following solves for n? n ≤ the quantity 2 times c minus 2 all over negative p n ≤ the quantity 2 times c minus 10 all over negative p n ≥ the quantity 2 times c minus 2 all over negative p n ≥ the quantity 2 times c minus 10 all over negative p

Gucchi:

am i right with C?

Gucchi:

i dont really know tbh

Gucchi:

@tetsxpreme

Gucchi:

does it matter if the p's are all negative or not?

Gucchi:

because i saw some questions online and the p's didnt have the negative

TETSXPREME:

The P is positive im pretty sure

Gucchi:

ohh...

Gucchi:

so that means im wrong

Gucchi:

maybe

TETSXPREME:

tbh idrk this one

Gucchi:

samee, i rlly just guessed

Gucchi:

@ihelpuuhelpme

Gucchi:

aiii @tranquility have any idea what the answer might be

Tranquility:

The first thing I would do is to distribute the 2 on the left side and then add 4 on both sides. What do you get once you've done that?

Tranquility:

When you're dealing with inequalities, you can treat it as an equal sign except when you're dividing by a negative number, that's when the sign flips But otherwise, the sign just remains the same so think of it as an equal sign for now −np − 4 = 2(c − 3) When you distribute 2 into (c-3), what do you get? a(b-c) = ab - ac

Gucchi:

so 2 times c which is 2c

Tranquility:

and what is 2 times 3

Gucchi:

6

Gucchi:

2c - 6?

Tranquility:

so 2(c - 3) is equal to 2*c - 2*3 Yes!

Tranquility:

We now have −np − 4 = 2c - 6 What do you get when you add 4 to both sides?

Gucchi:

- 10 = 2c - 10

Gucchi:

ohh wait

Gucchi:

just 2c - 10

Gucchi:

mb

Gucchi:

so i now know it could be the last option

Gucchi:

?

Tranquility:

no, you're doing 4 to -6 

Gucchi:

ohhh so its not - 10?

Tranquility:

No... what is -6 + 4 = ? You can even use a calculator

Gucchi:

oh then c right

Gucchi:

hold on

Gucchi:

-2

Tranquility:

We haven't gotten that far yet

Gucchi:

wait

Tranquility:

@gucchi wrote:
-2
Yes

Gucchi:

i mean b and d

Gucchi:

can be eliminated

Tranquility:

You confused me there for a second but yes B and D

Gucchi:

ok so now...?

Tranquility:

Now remember when I said this when you're dividing by a negative number, that's when the sign flips so > would become < and < would become >

Tranquility:

Right now we have −np ≤ 2c -2 We want to divide both sides by negative p since we're dividing by something negative, that means you have to flip the sign

Gucchi:

oh easy the final answer is C

Tranquility:

You got it

Gucchi:

ok so it really doesnt matter is the p is negative or not in this problem

Gucchi:

whether the p*

Tranquility:

P is just a variable We divided it by -1*p

Gucchi:

ohh alr makes sense

Tranquility:

Just dividing it by p would not have flipped the sign. You would then be left with -n and we want to get n all by itself. So you would have to divide by -1 and flip the sign

Gucchi:

thanks so much btw

Tranquility:

You're welcome!